OpenStudy (anonymous):

find critical pts. of sin^3(x^3)

OpenStudy (anonymous):

Find the derivative. Set the derivative equal to zero. Solve for x

OpenStudy (anonymous):

ugh. connection.

OpenStudy (anonymous):

hang on.

OpenStudy (anonymous):

\[\sin^3(x^3)\] let u equal x^3 \[\sin^3(u)\] \[\frac{ d }{ du } = \sin^3(u) = 3\sin^2(u)*\cos(u)\] Chain rule says \[f'(g(x))*g'(x); g(x)= x^3; f(x)= \sin^3(u)\] \[\frac{ d }{ dx }\sin^3(x^3) = 9x^2\sin^2(x^3)\cos(x^3)\]

OpenStudy (anonymous):

Oh, I thought it was just asking for the critical points

OpenStudy (anonymous):

Yeah, finding the maximums and minimums would be terrible, but we know what the critical points are.

OpenStudy (anonymous):

maxima and minima... >.<

OpenStudy (anonymous):

?